OpenStudy (anonymous):

I've been up all night doing homework and im either really stupid or forget how to do this. heeeeelp??? show the ratio 12 : 400 as a fraction in simplified form

ganeshie8 (ganeshie8):

\(a : b\) is same as \(\large \frac{a}{b}\)

ganeshie8 (ganeshie8):

\(12 : 400\) is same as \(\large \frac{12}{400}\) simplify...

OpenStudy (anonymous):

divide by two, right?

ganeshie8 (ganeshie8):

yes.. .looks \(4\) goes in both top and bottom

ganeshie8 (ganeshie8):

\(\large \frac{12}{400} \) \(\large \frac{4\times 3}{4 \times 100} \)

ganeshie8 (ganeshie8):

cancel 4

OpenStudy (anonymous):

oooooooooooooh thanks c:

ganeshie8 (ganeshie8):

np :) just making sure, final simplified form is what ? :)

OpenStudy (anonymous):

I dunno how to do fractions here but its 3 over 100 c:

ganeshie8 (ganeshie8):

thats right ! we simply cancel 4 as its common in both top and bottom. so final simplified form is 3/100